OpenStudy (anonymous):

Find the coordinates of the points of intersection of the graphs of x^2 - y^2 = 3, xy = 2, and y = -2x + 5.

OpenStudy (anonymous):

you have three equations and only two variables. maybe they do not intersect at all

OpenStudy (anonymous):

if \[x^2 - y^2 = 3\] and\[y=-2x+5\] then you can start by solving \[x^2-(-2x+5)^2=3\]

OpenStudy (anonymous):

once you get your two answers when you solve the quadratic, find of which pair also gives \(xy=2\)

OpenStudy (anonymous):

i get \(x=2\) or \(x=\frac{14}{3}\) from the quadratic equation